ccounting Rate of Return WeCare Clinic is planning on investing in some new echocardiogram equipment that...

  1. ccounting Rate of Return

    WeCare Clinic is planning on investing in some new echocardiogram equipment that will require an initial outlay of $175,000. The system has an expected life of five years and no expected salvage value. The investment is expected to produce the following net cash flows over its life: $89,000, $78,000, $92,000, $88,000, and $93,000.

    Required:

    1. Calculate the annual net income for each of the five years.

    Net Income
    Year 1 $
    Year 2 $
    Year 3 $
    Year 4 $
    Year 5 $

    2. Calculate the accounting rate of return. Enter your answer as a whole percentage value (for example, 16% should be entered as "16").
    %

    3. What if a second competing revenue-producing investment has the same initial outlay and salvage value but the following cash flows (in chronological sequence): $93,000, $93,000, $93,000, $89,000, and $36,000? Calculate its accounting rate of return. Enter your answer as a whole percentage value (for example, 16% should be entered as "16").
    %

    Using the accounting rate of return metric, which project should be selected: the first or the second?

    Why might the second project be preferred over the first project?
Answer #1
  1. Calculate net income:

initial outlay per year =175000/5

                                  = 35000

                 

Net income

Year 1

89000-35000 = 54000

Year 2

78000-35000 = 43000

Year 3

92000-35000 = 57000

Year 4

88000-35000 = 53000

Year 5

93000-35000 = 58000

2) Accounting rate of return = Average net income*100/Average investment

Average net income = 265000/5 = 53000

Accounting rate of return = 53000/175000= 0.30

3) Calculate net income:

Net income

Year 1

93000-35000 = 58000

Year 2

93000-35000 = 58000

Year 3

93000-35000 = 58000

Year 4

89000-35000 = 54000

Year 5

36000-35000 = 1000

Accounting rate of return = Average net income*100/Average investment

Average net income = 229000/5 = 45800

Accounting rate of return = 45800/175000 =0.26

as per the accounting rate of return metric first project should be selected because its accounting rate of return is higher than second one. first project is better off because it has strong rate of return.